# A Large Language Model Evolutionary Algorithm for Automatically Generating Metaheuristics

### v1

@@ Abstract
Large Language Models (LLMs) such as GPT-4 have demonstrated their ability to understand natural language and generate complex code snippets. This paper introduces a novel Large Language Model Evolutionary Algorithm (LLaMEA) framework, leveraging GPT models for the automated generation and refinement of algorithms. Given a set of criteria and a task definition (the search space), LLaMEA iteratively generates, mutates and selects algorithms based on performance metrics and feedback from runtime evaluations. This framework offers a unique approach to generating optimized algorithms without requiring extensive prior expertise. We show how this framework can be used to generate novel black-box metaheuristic optimization algorithms automatically. LLaMEA generates multiple algorithms that outperform state-of-the-art optimization algorithms (Covariance Matrix Adaptation Evolution Strategy and Differential Evolution) on the five dimensional black box optimization benchmark (BBOB). The results demonstrate the feasibility of the framework and identify future directions for automated generation and optimization of algorithms via LLMs.
@@ I Introduction
In this paper, we propose to overcome such limitations by using Large Language Models (LLMs) for automatically and iteratively evolving and optimizing the program code of such metaphor-based optimization algorithms. To test the approach, we use a specific optimization benchmarking tool to automatically evaluate the quality of the generated optimization algorithm and to provide a corresponding feedback to the LLM. In particular, the contributions of this work are as follows:
@@
  • We demonstrate that LLaMEA can generate a new metaphor-based optimization algorithm that successfully competes with the state-of-the-art algorithms on a standard set of benchmark test problems.
@@ II Related Work
Overall, these advancements illustrate the diverse use of LLMs in an optimization context. However, we observe that the application of LLMs for generating novel direct global optimization algorithms from scratch is still at a very early stage. Often, the performance evaluation of the LLM (e.g., in EvoLLM) or the generated algorithms (e.g., in AEL) focuses on small test problems or does not benchmark against state-of-the-art algorithms. In-context learning, as in EvoLLM, and algorithm generation, as in STOP or AEL, are usually not combined, and the generation of new algorithms is typically based on mutation and crossover requests to the LLM.
@@ III LLaMEA
9:         $F\leftarrow(S,((\text{name}(a_{0}),y_{0},\sigma_{0}),\ldots,(\text{name}(a_{t}),y_{t},\sigma_{t})),a_{b},e_{b},F_{0})$
@@
11:         $F\leftarrow(S,((\text{name}(a_{0}),y_{0},\sigma_{0}),\ldots,(\text{name}(a_{t}),y_{t},\sigma_{t})),a_{t},e_{t},F_{0})$
@@
The while-loop (line 7-20) iterates through $B$ (total budget) LLM calls to generate a new program each time (line 13). The prompt construction in lines 9 (in case of elitist selection) or line 11 (in case of non-elitist selection) is the essential step, generating a prompt $F$ that consist of a concatenation [^1] We use tuple notation here for clarity to show the components that compose the prompt string. of the following information: (i) Task description $S$; (ii) the algorithm names and mean fitness values and standard deviations of all previously generated algorithms; (iii) the current best algorithm’s code $a_{b}$ (in case of elitist selection) or the most recent algorithm’s code $a_{t}$ (in case of non-elitist selection); (iv) its execution error information $e_{t}$ (which is empty if there are no execution errors) and (v) a short task prompt $F_{0}$, with an instruction to the LLM. The new algorithm is generated by the LLM (line 13) and its fitness evaluated catching any runtime errors on the way (line 14), setting it explicitly to zero (worst possible fitness) if runtime errors occurred (line 15). The best-so-far algorithm is then updated if an improvement was found (line 17).
@@ III-A Starting Prompt
The initialisation of the optimization loop is crucial for the framework to work, as it sets the boundaries and rules that the LLM needs to operate with. Through experimentation, we found that including a small example code helps a lot in generating code without syntax and runtime errors. However, the example code could also bias the optimization search towards similar algorithms. In this work we therefore choose to use a simple implementation of random search as the example code to provide. The code of this random search algorithm is provided in our Zenodo repository [30]. The LLM receives an initial prompt with a specific set of criteria, domain expertise, and problem description. This starting point guides the LLM in generating an appropriate algorithm. The starting prompt has two main components, the specification of the role and general task e.g. ”Your task is to design novel metaheuristic algorithms to solve black box optimization problems.” and a detailed description of the problem to solve including a clear format of the expected response. An example problem description is given below:
@@ III-C Evaluation
The generated algorithm is evaluated on the Black-Box Optimization Benchmark (BBOB) suite [24]. The suite consists of $24$ noiseless functions, with an instance generation mechanic to generate a diverse set of different optimization landscapes that share particular function characteristics. The suite is divided into $5$ function groups; separable functions, functions with low or moderate conditioning, high conditioning unimodal functions, multi-modal functions with strong global structure and multi-modal functions with weak global structure. The BBOB suite is considered a best-practise in the field of metaheuristic algorithm design as evaluation of newly proposed algorithms. For the evaluation feedback provided to the LLM, we summarize the performance on the whole BBOB suite by taking an average any-time performance metric, in our case the area over the convergence curve (AOCC) (see Equation 1). This results in one number $y_{t}$ representing the performance of the proposed algorithm on average over the complete benchmark suite, and its standard deviation, $\sigma_{t}$. In addition to aggregating these values over all functions, we also experiment with a more detailed feedback mechanism where the average AOCC and its standard deviation are returned for each function group, resulting in ten performance values $(y_{t,1},\ldots,y_{t,5}),(\sigma_{t,1},\ldots,\sigma_{t,5})$. Theoretically, this approach should provide the LLM with information on what kind of functions the proposed solutions works well and on which it works less well.
@@ III-D Mutation, Selection and Feedback
After the selection is made, a feedback prompt $F$ is constructed using the following template:
@@
[…]
@@
We can ask the LLM to either make a (small) refinement of an algorithm or redesign it completely, where the latter is analogue to a restart or very large mutation rate in an ES optimization run, increasing its exploration behaviour. In the proposed framework we leave this choice to the LLM itself as our task is simply ”Either refine or redesign to improve the algorithm”. In Section V-B, we analyze how the LLM makes this decision over time, in terms of the generated names as well as the code similarity .
@@
Instead of providing an overall score and standard deviation, we can provide the LLM with more performance details of the generated algorithms. In our case, we can provide additional metrics per BBOB function group, potentially giving the LLM information on what kind of functions the solution works well and on which ones it does not work well. This results in mean AOCC values and standard deviations for each of the five function groups, i.e., $(y_{t,1},\ldots,y_{t,5}),(\sigma_{t,1},\ldots,\sigma_{t,5})$. In our experiments we cover the inclusion of such details versus leaving this information out. When we provided more detailed feedback, we provide the mean AOCC and the standard deviation of the AOCC scores for the separable functions, functions with low or moderate conditioning, unimodal functions with high conditioning, multi-modal functions with adequate global structure and functions with weak global structure (the five BBOB function groups). In practise, it would even be possible to provide the full convergence plots, either as csv or image to the multi-modal LLMs, this option we leave for future research.
@@
Instead of providing all codes, we keep a condensed list of algorithm names (generated by the LLM) and their respective scores and standard deviations of scores, $((\text{name}(a_{0}),y_{0},\sigma_{0}),\ldots,(\text{name}(a_{t}),y_{t},\sigma_{t}))$, to facilitate in-context learning of the LLM [25, 26]. The purpose of keeping this list and providing it to the LLM is two-fold, namely 1) the LLM could learn what kind of algorithms work well and which work less good (by analysing the algorithm names only), and 2) it makes it less likely that the LLM is generating the same algorithm twice.
@@ IV-B Benchmark Problems
To evaluate the proposed optimization algorithms we use the well-known Black-Box Optimization Benchmark suite (BBOB) [24] (See Table I in the Appendix for an overview of the functions). These benchmark functions can be separated into five groups with different high-level function characteristics. For a robust evaluation we use three different instances per function, where an instance of a BBOB function is defined by a series of random transformations that do not alter the global function characteristics. In addition, we perform $3$ independent runs per function instance with different random seeds (giving a total of $9$ runs per BBOB function). Each run has an evaluation budget of 10 0001000010\,00010 000 function evaluations. In our experiments we set the dimensionality of the optimization problems to $d=5$. We run the main (1 $\stackrel{{\scriptstyle+}}{{,}}\,$1)-LLaMEA optimization loop in Algorithm 1 for $T=100$ iterations.
@@ IV-C Performance Metrics
We aggregate the AOCC scores of all benchmark functions by taking the mean. The final mean AOCC determines the selected next parent (in a $(1+1)$-strategy) and is given as feedback to the LLM in the next mutation step.
@@ V Results and Discussion
Figure 2: Mean convergence curves (best-so-far algorithm scores) over the $5$ different runs for each LLM and strategy. Shaded areas denote the standard deviation of the best-so-far. w/Details denotes that we use a feedback mechanism that provides not just the plain average AOCC but also the average AOCC per BBOB function group as feedback to the LLM.
@@ V-B Mutation Rates
To analyse how much the LLMs change (mutate) the algorithms over an entire optimization run, the code diff ratio (number of code lines that are different between a pair of programs divided by the length of the largest code) is calculated over each run between parent and offspring solutions. In Figure 4, the mean diff over $5$ different runs per LLM is shown. There are a few interesting observations we can make: The difference between parent and offspring for GPT-3.5 is on average much lower than for other models, the $(1+1)$-GPT3.5-LLaMEA shows much higher differences in the beginning of the run than the other strategies with the same model. GPT-4-LLaMEA in general shows the largest differences (exploration) over the runs. It is very interesting to observe that the ratio of code differences in most cases seems to converge, indicating more exploration in the beginning of the search and more exploitation during the final parts of the search. This is interesting as the LLM has no information on the search budget, and can therefore also not base its decision to make large or small refinements on the stage of the optimization run. We can furthermore see in the generated names of the algorithms that the LLM has either tried to refine the algorithm (adding “Improved” or “Refined” or “Enhanced” to the name, or names such as <algorithm>V1, etc.) or redesign the algorithm using a different strategy (based on different existing algorithm names such as Differential Evolution, Particle Swarm Optimization etc.). We can therefore also look at the similarity between parent and offspring algorithm names to visualize the refinement process of the optimization runs.
@@
Figure 5 illustrates the average algorithm name similarity over iterations for each optimization run and LLM. Figure 5 illustrates the Jaro similarity mean (five runs) of the subsequently generated algorithm names (i.e., $\text{Jaro}(\text{name}(a^{\prime}),\text{name}(a_{t+1}))$ for $t=0,\ldots,B-1$, where $a^{\prime}=a_{b}$ for $(1+1)$-selection and $a^{\prime}=a_{t}$ for $(1,1)$-selection) over iterations $t$ for each of the nine LLM-based optimization run configurations. In Figure 6, we show the Jaro similarity for just a single run selected for three of the nine configurations.
@@ VI Analysis of Best Algorithms
The experiment above resulted in 3 65736573\,6573 657 algorithms (out of a maximum [^2] Resulting from $5$ runs with $T=100$ iterations, for $9$ different models. of 4 50045004\,5004 500) that were at least able to get an AOCC score larger than zero on the BBOB suite of optimization problems in $5$ dimensions. In the next step to validate our proposed framework, we evaluate how much the best of these algorithms can generalize beyond the evaluation performed during the optimization loop, and subsequently we analyze the code and behaviour of the best-performing algorithms that beat the state-the-art baseline in-depth. The evaluation is done in $5$, $10$ and $20$ dimensions, using $5$ different instances and $5$ independent runs per instance (so $25$ runs per BBOB function to optimize). The state-of-the-art baselines include the covariance matrix adaptation evolution strategies (CMA-ES) [40], BIPOP-CMA-ES [41] and Differential Evolution (DE) [42].
@@
In Figure 7 the empirical attainment functions (EAF) [37] are shown for the best algorithm generated per LLM and strategy, so $9$ algorithms in total. The empirical attainment function estimates the percentage of runs that attain an arbitrary target value not later than a given runtime. Taking the partial integral of the EAF results in a more accurate version of the Empirical Cumulative Distribution Function, since it does not rely on discritization of the targets. The area under this curve is a similar measure as the AOCC used during the LLaMEA optimization runs, and denotes how well an optimization algorithm performed over the complete run (on average over all instances, independent runs and objective functions). From the EAF curves and the area under these curves (see Table II in the Appendix), we can observe that LLaMEA has found two algorithms with better performance for $d=5$ than the state-of-the-art CMA-ES, BIPOP-CMA-ES and DE, with the algorithm called RADEA being the top performer and ERADS_QuantumFluxUltraRefined the second best. We therefore analyze both RADEA and the second best algorithm, ERADS_QuantumFluxUltraRefined (ERADS) in detail in the next sections. Detailed convergence curves per BBOB function (for $d=5$) are shown in Figure 8.
@@
Figure 8: Median best-so-far function value ($y$-axis) over the 10 0001000010\,00010 000 function evaluations ($x$-axis) per BBOB function in $5d$ for the best algorithm per LLM configuration (9 algorithms) and the baselines; CMA-ES, DE and BIPOP-CMA-ES.
VI-A Robust Adaptive Differential Evolution with Archive
One of the two top performing algorithms generated by the LLaMEA is ”Robust Adaptive Differential Evolution with Archive” (RADEA). The key idea of RADEA is the utilization of an archive of best solutions, which are used in the mutation operator.
Algorithm 2 Robust Adaptive Differential Evolution with Archive (RADEA)
1:$N\leftarrow 20$triangleright Population size
2:$N^{\prime}\leftarrow 50$triangleright Archive size
3:$P\leftarrow$ u.a.r. population initialization within $[low,high]$
4:$A\leftarrow\emptyset$triangleright Initialize archive as empty
5:$t\leftarrow 0$triangleright Initialize iteration counter
6:$F\leftarrow 0.8,CR\leftarrow 0.9$triangleright Control parameters for mutation and crossover
7:$\text{fit}_{P}\leftarrow f(P)$triangleright Evaluate initial population
8:while $t\cdot N<B$ do $\triangleright$ While budget not exhausted
9:     $P^{\prime}\leftarrow$ Mutation$(P,A,F)$triangleright Generate mutants
10:     $C\leftarrow$ Crossover$(P,P^{\prime},CR)$triangleright Crossover to create trial population
11:     $\text{fit}_{C}\leftarrow f(C)$triangleright Evaluate trial population
12:     $P,\text{fit}_{P}\leftarrow$ Selection$(P,\text{fit}_{P},C,\text{fit}_{C})$triangleright Selection of the fittest
13:     Update_Archive$(A,N^{\prime},P,\text{fit}_{P})$triangleright Update archive with improved solutions
14:     $t\leftarrow t+1$triangleright Increment iteration counter
15:end while
16:$i_{best}\leftarrow\text{argmin}(\text{fit}_{P})$triangleright Index of the best solution
17:return $P[i_{best}],\text{fit}_{P}[i_{best]}$triangleright Return best solution and its fitness
The pseudocode of RADEA, representing the key ideas of the generated Python code, was extracted manually and is shown in Algorithm 2. The full Python code can be retrieved from our online repository [30]. The different functions of RADEA are summarized as follows:
  • Mutation$(P,A,F)$: Generates mutants for each individual in the population by selecting three distinct vectors from the union of $P$ and $A$ (if $A$ is not empty), creating a new vector by adding the weighted difference of two vectors to the third. This method ensures that the archive can contribute to the diversity of the mutation process.
  • Crossover$(P,P^{\prime},CR)$: Performs crossover between the population $P$ and the mutant population $P^{\prime}$ with a crossover rate $CR$. Each dimension of each individual has a probability $CR$ to be replaced by the corresponding dimension in the mutant.
  • Selection$(P,\text{fit}_{P},C,\text{fit}_{C})$: Compares each individual in $P$ with its counterpart in $C$. An individual in $C$ with better fitness replaces the corresponding individual in $P$.
  • Update_Archive$(A,N^{\prime},P,\text{fit}_{P})$: Adds improved solutions to the archive $A$ and ensures that the size of the archive does not exceed its maximum capacity $N^{\prime}$ by removing the oldest entries.
The proposed RADEA algorithm seems most similar to the existing JADE [43] algorithm from literature. There are however some key differences which are, to the best of our knowledge, novel and have not been published before.
The key differences primarily revolve around the use of archives and parameter adaptation. RADEA emphasizes a robust use of an archive not only to store but actively use archived solutions in the mutation process, enhancing diversity and potentially aiding in overcoming local optima. This archive dynamically influences mutation strategies and contributes directly to the generation of new candidate solutions. In contrast, JADE’s use of an archive is optional and primarily serves to maintain diversity without a direct integration into the mutation process. Furthermore, JADE features adaptive control parameters (F and CR) that adjust dynamically based on the success rates of previous generations, aiming to optimize these parameters in real time to improve performance. RADEA does not show adaptive control of parameters even though the name suggests otherwise.
VI-B ERADS Quantum Flux Ultra Refined
@@
While the name sounds rather futuristic and very sophisticated, upon a close inspection of the code the ”ERADS Quantum Flux Ultra Refined” algorithm looks very similar to the previously described RADEA algorithm, which is interesting as the RADEA algorithm was generated during a different optimization run (though with the same GPT model). This could indicate a bias towards Differential Evolution style algorithms in the search. According to the LLM, ERADS stands for ”Enhanced RADEDM with Strategic Mutation”, and RADEDM stands for ”Refined Adaptive Differential Evolution with Dynamic Memory”. The pseudocode of ERADS was again extracted manually and is shown in Algorithm 3. The full Python code can be retrieved from our online repository [30].

### v2

@@ Abstract
Large Language Models (LLMs) such as GPT-4 have demonstrated their ability to understand natural language and generate complex code snippets. This paper introduces a novel Large Language Model Evolutionary Algorithm (LLaMEA) framework, leveraging GPT models for the automated generation and refinement of algorithms. Given a set of criteria and a task definition (the search space), LLaMEA iteratively generates, mutates and selects algorithms based on performance metrics and feedback from runtime evaluations. This framework offers a unique approach to generating optimized algorithms without requiring extensive prior expertise. We show how this framework can be used to generate novel black-box metaheuristic optimization algorithms automatically. LLaMEA generates multiple algorithms that outperform state-of-the-art optimization algorithms (Covariance Matrix Adaptation Evolution Strategy and Differential Evolution) on the five dimensional black box optimization benchmark (BBOB). The algorithms also show competitive performance on the 10- and 20-dimensional instances of the test functions, although they have not seen such instances during the automated generation process. The results demonstrate the feasibility of the framework and identify future directions for automated generation and optimization of algorithms via LLMs.
@@ I Introduction
In this paper, we propose to overcome such limitations by using Large Language Models (LLMs) for automatically and iteratively evolving and optimizing the program code of such metaphor-based optimization algorithms. To test the approach, we use a specific optimization benchmarking tool to automatically evaluate the quality of the generated optimization algorithm and to provide a corresponding feedback to the LLM. The specific contributions of this work are as follows:
@@
  • We demonstrate that LLaMEA can generate new metaphor-based optimization algorithms that successfully compete with state-of-the-art algorithms on a standard set of benchmark test problems.
@@ II Related Work
The recently proposed FunSearch approach (searching in the function space) [28] closes the loop and combines the LLM with a systematic evaluation of the quality of the generated programs, resulting in new solution algorithms for combinatorial optimization problems (cap-set problem, bin packing). In this approach, a distributed island-based Evolutionary Algorithm is used for maintaining diversity and prompt construction based on the already generated programs.
@@
Overall, these advancements illustrate the diverse use of LLMs in an optimization context. However, we observe that the application of LLMs for generating novel direct global optimization algorithms from scratch is still at a very early stage. Often, the performance evaluation of the LLM (e.g., in EvoLLM) or the generated algorithms (e.g., in AEL) focuses on small test problems or does not benchmark against state-of-the-art algorithms. In-context learning, as in EvoLLM, and algorithm generation, as in STOP or AEL, are usually not combined, and the generation of new algorithms is typically based on mutation and crossover requests to the LLM, with a focus on solving combinatorial problems.
@@ III LLaMEA
9:         $F\leftarrow(S,((\text{name}(a_{0}),y_{0}),\ldots,(\text{name}(a_{t}),y_{t})),(a_{b},y_{b},\sigma_{b},e_{b}),F_{0})$
@@
11:         $F\leftarrow(S,((\text{name}(a_{0}),y_{0}),\ldots,(\text{name}(a_{t}),y_{t})),(a_{t},y_{t},\sigma_{t},e_{t}),F_{0})$
@@
The while-loop (line 7-20) iterates through $T$ (total budget) LLM calls to generate a new program each time (line 13). The prompt construction in lines 9 (in case of elitist selection) or line 11 (in case of non-elitist selection) is the essential step, generating a prompt $F$ that consist of a concatenation [^1] We use tuple notation here for clarity to show the components that compose the prompt string. of the following information: (i) Task description $S$; (ii) the algorithm names and mean fitness values of all previously generated algorithms; (iii) the current best algorithm’s $a_{b}$ (in case of elitist selection) or most recent algorithm’s $a_{t}$ (in case of non-elitist selection) complete information (code, mean fitness and standard deviation, execution error information) and (iv) a short task prompt $F_{0}$, with an instruction to the LLM. The new algorithm is generated by the LLM (line 13) and its fitness evaluated catching any runtime errors on the way (line 14), setting $y_{t+1}$ explicitly to zero (worst possible fitness) if runtime errors occurred (line 15). The best-so-far algorithm is updated if an improvement was found (line 17).
@@ III-A Starting Prompt
The initialisation of the optimization loop is crucial for the framework to work, as it sets the boundaries and rules that the LLM needs to operate with. Through experimentation, we found that including a small example code helps a lot in generating code without syntax and runtime errors. However, the example code could also bias the optimization search towards similar algorithms. In this work we therefore choose to use a simple implementation of random search as the example code to provide. The code of this random search algorithm is provided in our Zenodo repository [31]. The LLM receives an initial prompt with a specific set of criteria, domain expertise, and problem description. This starting point guides the LLM in generating an appropriate algorithm. The starting prompt has two main components, the specification of the role and general task, e.g., ”Your task is to design novel metaheuristic algorithms to solve black box optimization problems.”, and a detailed description of the problem to solve including a clear format of the expected response. Our detailed task prompt $S$ is given below:
@@ III-C Evaluation
The generated algorithm is evaluated on the Black-Box Optimization Benchmark (BBOB) suite [24]. The suite consists of $24$ noiseless functions, with an instance generation mechanic to generate a diverse set of different optimization landscapes that share particular function characteristics. The suite is divided into $5$ function groups; separable functions, functions with low or moderate conditioning, high conditioning unimodal functions, multi-modal functions with strong global structure and multi-modal functions with weak global structure. The BBOB suite is considered a best-practice in the field of metaheuristic algorithm design as evaluation of newly proposed algorithms. For the evaluation feedback provided to the LLM, we summarize the performance on the whole BBOB suite by taking an average any-time performance metric, in our case the area over the convergence curve (AOCC) (see Equation 1). This results in one number $y_{t}$ representing the performance of the proposed algorithm on average over the complete benchmark suite, and its standard deviation, $\sigma_{t}$. In addition to aggregating these values over all functions, we also experiment with a more detailed feedback mechanism where the average AOCC and its standard deviation are returned for each function group, resulting in ten performance values $(y_{t,1},\ldots,y_{t,5}),(\sigma_{t,1},\ldots,\sigma_{t,5})$. Theoretically, this approach should provide the LLM with information on what kind of functions the proposed solutions works well and on which it works less well.
@@ III-D Mutation, Selection and Feedback
After the selection is made, a feedback prompt $F$ (lines 9, 11 of Algorithm 1) is constructed using the following template:
@@
The prompt includes the initial detailed task prompt $S$, a list $((\text{name}(a_{0}),y_{0}),\ldots,(\text{name}(a_{t}),y_{t}))$ of previously generated algorithm names and their mean scores, the selected algorithm $a_{b}$ or $a_{t}$, including its score $y_{b}$ ($y_{t}$) and standard deviation $\sigma_{b}$ ($\sigma_{t}$), to mutate and a short task prompt $F_{0}=$ ”Either refine or redesign to improve the algorithm” telling the LLM to perform a mutation or redesign (restart) action. The list of previously tried algorithm names is included to make sure the LLM is not generating (almost) the same algorithm twice.
@@
We ask the LLM to either make a (small) refinement of an algorithm or redesign it completely, where the latter is analogue to a restart or very large mutation rate in an ES optimization run, increasing its exploration behaviour. In the proposed framework we leave this choice to the LLM itself as our task is simply ”Either refine or redesign to improve the algorithm”. In Section V-B, we analyze how the LLM makes this decision over time, in terms of the generated names as well as the code similarity .
@@
Instead of providing an overall score and standard deviation, we can provide the LLM with more performance details of the generated algorithms. In our case, we can provide additional metrics per BBOB function group, potentially giving the LLM information on what kind of functions the solution works well and on which ones it does not work well. This results in mean AOCC values and standard deviations for each of the five function groups, i.e., $(y_{t,1},\ldots,y_{t,5}),(\sigma_{t,1},\ldots,\sigma_{t,5})$. In our experiments we cover the inclusion of such details versus leaving this information out. When we provided more detailed feedback, we provide the mean AOCC and the standard deviation of the AOCC scores for the separable functions, functions with low or moderate conditioning, unimodal functions with high conditioning, multi-modal functions with adequate global structure and functions with weak global structure (the five BBOB function groups). In practise, it would even be possible to provide the full convergence plots, either as csv or image to the multi-modal LLMs; we leave this option for future research.
@@
Instead of providing all codes, we keep a condensed list of algorithm names (generated by the LLM) and their respective scores, $((\text{name}(a_{0}),y_{0}),\ldots,(\text{name}(a_{t}),y_{t}))$, to facilitate in-context learning of the LLM [25, 26]. The purpose of keeping this list and providing it to the LLM is two-fold, namely 1) the LLM could learn what kind of algorithms work well and which work less good (by analysing the algorithm names only), and 2) it makes it less likely that the LLM is generating the same algorithm twice.
@@ IV-B Benchmark Problems
[…]
@@ IV-C Performance Metrics
We aggregate the AOCC scores of all benchmark functions by taking the mean. The final mean AOCC is given as feedback to the LLM in the next step, and is used as best-so-far solution in case an improvement was found.
@@ V Results and Discussion
Figure 2: Mean convergence curves (best-so-far algorithm scores) over the $5$ different runs for each LLM and strategy. Shaded areas denote the standard deviation of the best-so-far. w/Details denotes that we use a feedback mechanism that provides not just the plain average AOCC (and its standard deviation) but also the average AOCC and standard deviation per BBOB function group as feedback to the LLM.
@@ V-B Mutation Rates
To analyse how much the LLMs change (mutate) the algorithms over an entire optimization run, the code diff ratio (number of code lines that are different between a pair of programs divided by the length of the largest code) is calculated over each run between parent and offspring solutions. In Figure 4, the mean diff over $5$ different runs per LLM is shown.
There are a few interesting observations we can make: The difference between parent and offspring for GPT-3.5 is on average much smaller than for other models, the $(1+1)$-GPT3.5-LLaMEA shows much higher differences in the beginning of the run than the other strategies with the same model. GPT-4-LLaMEA in general shows the largest differences (exploration) over the runs. It is very interesting to observe that the ratio of code differences in most cases seems to converge, indicating more exploration in the beginning of the search and more exploitation during the final parts of the search. This is interesting as the LLM has no information on the search budget $T$ of Algorithm 1, and can therefore also not base its decision to make large or small refinements on the stage of the optimization run. We can furthermore see in the generated names of the algorithms that the LLM has either tried to refine the algorithm (adding “Improved” or “Refined” or “Enhanced” to the name, or names such as <algorithm>V1, etc.) or redesign the algorithm using a different strategy (based on different existing algorithm names such as Differential Evolution, Particle Swarm Optimization etc.). We can therefore also look at the similarity between parent and offspring algorithm names to visualize the refinement process of the optimization runs.
@@
Figure 5 illustrates the Jaro similarity mean (five runs) of the subsequently generated algorithm names (i.e., $\text{Jaro}(\text{name}(a^{\prime}),\text{name}(a_{t+1}))$ for $t=0,\ldots,B-1$, where $a^{\prime}=a_{b}$ for $(1+1)$-selection and $a^{\prime}=a_{t}$ for $(1,1)$-selection) over iterations $t$ for each of the nine LLM-based optimization run configurations. In Figure 6, we show the Jaro similarity for just a single run selected for three of the nine configurations.
@@ VI Analysis of Best Algorithms
The experiment above resulted in 3 65736573\,6573 657 algorithms (out of a maximum2225555 runs with $T=100$ iterations, each, for $9$ different models. of 4 50045004\,5004 500) that were at least able to get an AOCC score larger than zero on the BBOB suite of optimization problems in $5$ dimensions. In the next step to validate our proposed framework, we evaluate how much the best of these algorithms can generalize beyond the evaluation performed during the optimization loop, and subsequently we analyze in-depth the code and behaviour of the best-performing algorithm that beats the state-the-art baselines. The evaluation is done in $5$, $10$ and $20$ dimensions, using $5$ different instances and $5$ independent runs per instance (so $25$ runs per BBOB function to optimize). For a fair comparison, we decided to re-run the generated algorithms for $d=5$ in this setting (with $25$ runs), too. The state-of-the-art baselines are the covariance matrix adaptation evolution strategies (CMA-ES) [41], Differential Evolution (DE) [42] and an extensively optimised version of Modular CMA-ES [37] denoted as CMA-best. The optimised CMA-ES algorithm is the best-configured algorithm (in terms of AOCC) on BBOB in $5$ dimensions over more than 52 1285212852\,12852 128 CMA-ES and DE algorithm variants, note that beating this optimized baseline is incredibly hard as it was the result of a very large modular optimization experiment. We use the CMA-best baseline to visualise a reasonable upper-bound of what is possible when optimizing an algorithm configuration to a specific benchmark set and dimensionality.
@@
In Figure 7 the empirical attainment functions (EAF) [38] are shown for the best algorithm generated per LLM and strategy, so $9$ algorithms in total plus the $3$ baselines. The empirical attainment function estimates the percentage of runs that attain an arbitrary target value not later than a given runtime. Taking the partial integral of the EAF results in a more accurate version of the Empirical Cumulative Distribution Function, since it does not rely on discritisation of the targets. The area under this curve is a similar measure as the AOCC used during the LLaMEA optimization runs, and denotes how well an optimization algorithm performed over the complete run (on average over all instances, independent runs and objective functions). From the EAF curves and the area under these curves (see Table II in the Appendix), we can observe that LLaMEA has found one algorithms (ERADS_QuantumFluxUltraRefined, for short ERADS) with better performance for $d=5$ than the state-of-the-art CMA-ES in terms of AOCC, two algorithms (ERADS and AdaptiveDifferentialEvolutionHarmonySearch) that on average perform better than the optimized CMA-best after the total budget in $5d$ (the EAF curve reaches higher) and six algorithms that perform better than the DE baseline but less than the CMA baseline Detailed convergence curves per BBOB function (for $d=5$) are shown in Figure 8. From this Figure we can observe that especially for BBOB $f_{17}$ and $f_{18}$, the ERADS algorithm shows very promising search behaviour.
@@
Figure 8: Median best-so-far function value ($y$-axis) over the 10 0001000010\,00010 000 function evaluations ($x$-axis) per BBOB function in $5d$ for the best algorithm per LLM configuration (9 algorithms) and the baselines; CMA-ES, DE and CMA-best.
ERADS Quantum Flux Ultra Refined
@@
While the name sounds rather futuristic and very sophisticated, upon a close inspection of the code the ”ERADS Quantum Flux Ultra Refined” algorithm looks very similar a standard Differential Evolution algorithm. According to the LLM, ERADS stands for ”Enhanced RADEDM with Strategic Mutation”, and RADEDM stands for ”Refined Adaptive Differential Evolution with Dynamic Memory”. The pseudocode of ERADS, representing the key ideas of the generated Python code, was extracted manually and is shown in Algorithm 2. The full Python code can be retrieved from our online repository [31].